OpenStudy (anonymous):

I need help with trig identities! I can I prove this as a trig identity? 2sec^2(x)-2tan^2(x)/csc(x) = sin2(x)sex(x) HELP PLEASE

OpenStudy (mathmale):

I'm a little worried about that " sex(x) ." Care to clarify your meaning? :)

OpenStudy (anonymous):

i LOL SORYY! i meant sec(x)

OpenStudy (mathmale):

2sec^2(x)-2tan^2(x)/csc(x) = sin2(x)sec(x) Could be written as 2 (sec x)^2 - [2(tan x)^2]/(csc x) =(sin x)^2 sec x. Is this also your interpretation?

OpenStudy (anonymous):

Yeah i think you can write it like that too!

OpenStudy (mathmale):

Next, you might want to re-write your (sec x)^2 term as 1 / (cos x)^2. Using the same approach, rewrite (tan x)^2 as (sin x)^2 / (cos x)^2. finally, express your sec x in the right side as 1 / cos x. Often it's easier to do proofs when we're dealing only with the basic functions sine and cosine.

OpenStudy (anonymous):

i'm still a bit confused... here's the question...